# Deep Chandra X-ray Observations of Abell 2029: the Merger History of a Relaxed, Strong Cool Core Cluster

### v1

@@ 4.2 Temperature and Abundance Maps
Spectral maps of A2029 were created following the methods of S. Randall et al. (2008); S. W. Randall et al. (2009). For each spectral map pixel (98 pixel-1), spectra were extracted from circular regions whose radius was allowed to increase until a minimum of 20,000 background-subtracted counts were contained. The radius of the extraction region varies from 10″ ( 14 kpc) in bright regions near the core to ( 509 kpc) in faint regions near the edges of the map. Within this region, the spectrum of each ObsID is extracted and then fit simultaneously using an absorbed APEC model, with only the hydrogen column density — — and redshift — — frozen during fitting.
@@
The innermost spiral excess (highlighted by the white contours) traces keV gas temperatures and higher metallicities, as expected from sloshing ICM. The SW “splash” feature (white arrows in Fig. 7) traces regions of keV gas and leads into the coldest regions, with temperatures around keV. This cold region was previously seen in XMM and Chandra observations of A2029 (H. Bourdin & P. Mazzotta, 2008; S. A. Walker et al., 2012). The splash also leads into a region of higher ( Z⊙) metallicity (Fig. 8). The region of isolated excess emission to the SW of the cluster (green arrow in Fig. 7) corresponds to a region of colder ( keV) gas, which may be associated with a background group or cluster (discussed further in §5.5). A large ( 0.9 Mpc) hot ( keV) region to the NW of the cluster center may indicate shock-heated gas in the region.
@@ 5.2 Potential Merger Shock
We used the python package pyproffit (v0.7.0; D. Eckert et al., 2020) to plot the surface brightness profile in a region over the large hot region observed to the NW in the temperature map (Fig. 7). We used a background subtracted image with point sources removed, an exposure map in units of seconds, and a user-generated error map to extract the surface brightness profile of the shock region using pyproffit. The surface brightness profile was extracted in a circular sector spanning 1.5∘ to , centered on the peak of the X-ray emission, with a maximum radial extent of 630″, and a bin size of 10″ in order to ensure a minimum of 5,000 net counts in the outermost bin.
We fit the 2D surface brightness within the kpc [300″] to kpc [630″] range using a projected broken power-law 3D density model (M. Markevitch & A. Vikhlinin, 2007; M. S. Owers et al., 2009; J. Tiwari & K. P. Singh, 2022; K. Rajpurohit et al., 2024) of the form:
@@
We can determine the Mach number of the potential shock front using a similar equation as Eq. 2, but using the temperature jump, expressed as a function of the density jump (J. S. Sanders et al., 2016b):
where and are the upstream and downstream gas temperatures, respectively. For regions where , when the temperature ratio, , is greater than unity it suggests the presence of a shock front, while ratios of less than unity indicate a potential cold front (K. Rajpurohit et al., 2024).
@@ 5.3.1 Bay Feature and SW Excess
As noted in §3, there appears to be a concave bay-like structure located kpc south of the cluster core that could be attributed to a Kelvin-Helmholtz instability (KHI). Similar structures have been reported in previous clusters (e.g., Perseus, Centaurus, and Abell 1795; J. S. Sanders et al., 2016a; S. A. Walker et al., 2014, 2017) and are thought to be associated with large-scale KHI bays.
@@ 5.3.2 SE ‘Splash’
In §3, we noted a region of excess emission SE of the cluster core, which we dubbed the “splash” feature, noted in Fig. 3 with white arrows. Figure 15 illustrates residual emission (left panel) overlaid with the annular sectors used to extract the azimuthal surface brightness profiles (right panel). We compared the SE direction (solid squares) with the northwestern (NW) direction (open circles) in the right panel. The SW splash excess is observed between 196∘ to 247∘, highlighted in purple in Fig. 15. The source of this extended feature could be due to the second, perturbing cluster system. Some simulations (J. A. ZuHone et al., 2010; J. A. ZuHone & E. Roediger, 2016) have shown similar “splash”-like features, in conjunction with sloshing spiral structures, that results after the second core passage of the perturbing body.
@@ 5.3.3 AGN Bubbles and Feedback
At the core of A2029 lies the AGN radio source PKS1508+059. AGN are known to inflate bubbles in the ICM that can offset radiative cooling. Previous studies of A2029 have been unable to detect any significant555D. A. Rafferty et al. (2006) report the detection of one cavity, but at low significance. evidence of such bubbles in X-ray imaging. While the observations here similarly reveal no clear evidence for such bubbles, we also cannot rule out the possibility that the southern bay feature could be associated with the inner edge of an outer ghost cavity that could have been inflated from the southern lobe. A similar situation arose in the Ophiuchus cluster, where a concave X-ray edge was initially deemed unrelated to the radio emission, but much deeper low-frequency radio data later revealed a matching ghost cavity (S. Giacintucci et al., 2020).
Figure 16 shows the X-ray and residual emission images in the inner 5′ 5′ cluster core to highlight the central WAT source. The overlaid cyan contours are the 1490 MHz radio emission taken from VLA observations presented in G. B. Taylor et al. (1994). Contours are drawn at levels corresponding to 0.0006, 0.0024, 0.0079, and 0.017 Jy beam-1. The X-ray images show no obvious signs of isolated regions of decreased emission near the radio lobes, nor anywhere in the ICM at larger radii. In the -model subtracted residual emission image, though showing decreased emission in the regions of both lobes, the observed depressions appear to be more closely associated with the cooling spiral structure rather than with distinct cavities. However, cavities may be particularly hard to detect in these regions because of confusion with the sloshing features.
@@
Recent work by R. Timmerman et al. (2022) applied a hybrid X-ray–radio method to estimate AGN cavity powers in galaxy clusters. For A2029, they utilize high-resolution () observations from the International Low Frequency Array (LOFAR; M. P. Van Haarlem et al., 2013) telescope (ILT) and low-resolution imaging from the LOFAR Two-Metre Sky Survey (LoTSS; T. W. Shimwell et al., 2017, 2019, 2022). While the authors report estimates of cavity volumes and power, they do not identify definitive X-ray cavities coincident with the radio lobes, noting that the complex spiral morphology of the ICM complicates the clear identification of such features.
We found in §4 a mass deposition rate within the central 40″ (equivalent to 58 kpc at the redshift of A2029) of M⊙ yr-1 when the low temperature of the cooling flow model is free. However, we find when the low temperature is fixed to 0.0808 keV and the high temperature is tied to an APEC component. The latter is the more relevant quantity to compare to the star formation rate of the cD galaxy — A. K. Hicks et al. (2010) measure a SFR of only M⊙ yr-1 and A. Fraser-McKelvie et al. (2014) estimate M⊙ yr-1. The large discrepancy in the results between the free- mass deposition rate and the SFR, coupled with the much smaller fixed- mass deposition rate, indicates that in A2029, most gas does not cool to sub-keV temperatures, consistent with a hidden or partial cooling flow scenario (A. C. Fabian et al., 2022).
To see if the central AGN is providing the heating necessary to offset the observed cooling, we compared the energy injection rate, , from the radio source to the luminosity of the cooling gas, . For A2029, R. Paterno-Mahler et al. (2013) estimated an energy injection rate666Calculated as , where and are the cavity pressure and volume, respectively, averaged over a typical AGN repetition rate of yr. of erg s-1. Additionally, R. Timmerman et al. (2022) used low-resolution LOFAR observations to estimate a total cavity power (i.e., the AGN injection rate) of erg s-1; however the authors assumed radial propagation of the radio lobes, while it’s clear that for A2029, the lobes are bent relative to the jet axis.
@@ 5.4 Comparison to Simulation
We compare our observations to simulations of galaxy cluster mergers in order to get a better picture of the past merger history that would result in some of the features we have identified in A2029 so far. We examined the idealized binary merger simulations from the Galaxy Cluster Merger Catalog777https://gcmc.hub.yt (J. A. ZuHone et al., 2018) and found that our observations were qualitatively reproduced by a simulation of a 1:10 mass ratio merger, with an impact parameter of 500 kpc, originally presented in J. A. ZuHone (2011).
@@ 5.5 Optical Structure
J. Sohn et al. (2019) identify spectroscopic members of A2029, which we show overlaid on the X-ray image in the left panel of Fig. 20 in green circles. We create contours (in steps of 0.28 galaxies arcmin-1) of galaxy surface density, following J. Sohn et al. (2019), which we show as white contours in all panels of Fig. 20.
@@
In addition to member galaxies, J. Sohn et al. (2019) show that there are two infalling subsystems, consisting of Abell 2033 and a newly identified group, the Southern Infalling Group (SIG), based on weak lensing and X-ray observations, and phase-space analysis of the member galaxies. Both systems lie outside the Chandra FoV, which is roughly the size of the cluster radius Mpc (J. Sohn et al., 2019).
In the X-ray and residual emission images, we identified a region of excess emission located kpc SW of the cluster core (highlighted with green arrows in Figs. 3; see §3.1.1). In §5.3, we determined that this emission was not likely to be associated with the main cluster and could instead suggest the presence of a foreground or background group or cluster, with its own intragroup or intracluster medium that we are detecting. The thumbnail image of the radio source, shown in Fig. 2 panel (b), reveals the presence of numerous red sources clustered together. J. Sohn et al. (2019) measures a spectroscopic redshift of , attributed to the largest and brightest source of the group shown in Fig. 2 panel (b). The concurrence with radio emission indicates the presence of an AGN within the group of red galaxies. In addition, this group coincides with the region of excess X-ray emission seen to the SW of the cluster (green arrow in Fig. 3). All this together supports the picture of a background group or cluster, with its own intragroup or intracluster medium, that lies at a redshift .
The southernmost double-lobe radio source, located 680 kpc nearly due south of the cluster center, is at a redshift of (J. Sohn et al., 2019), indicating it could be associated with the SW radio source, which lies at similar redshift. The thumbnail of the source, shown in panel (c) of Fig. 2, shows two sources near the expected location of the radio source, with the one identified as a point source during the wavdetect run (see §3) indicated by a red circle. This source is likely the AGN driving the two radio lobes seen in 1.4 GHz.
@@ 6 Summary and Conclusions
We find a global, best-fit projected temperature of keV and abundance of within the inner 116″[ kpc]. This is slightly lower temperature than previous studies (e.g., R. Paterno-Mahler et al. (2013)) but consistent within the combined statistical and systematic uncertainties. A two-temperature model provides a significantly improved fit, with components at keV and keV, yielding much tighter constraints on the hot temperature component than earlier studies (T. E. Clarke et al., 2004a; R. Paterno-Mahler et al., 2013). Fitting a cooling flow model returned best-fit temperatures of keV and keV, abundance , and mass deposition rate M⊙ yr-1. This is in agreement with previous studies that have reported similarly high cooling flow rates (C. L. Sarazin & R. W. O’Connell, 1993; C. B. Peres et al., 1998; R. Paterno-Mahler et al., 2013), despite the relatively low levels of star formation in the central cD galaxy compared to expectations from unimpeded cooling (B. McNamara & P. Nulsen, 2007; M. McDonald et al., 2010; A. K. Hicks et al., 2010). The low temperature component is consistent with the latest XRISM findings which report a 3.42 keV temperature component detected within the inner kpc [2.5′] (A. Sarkar et al., 2025). However, when keeping the low temperature fixed to 0.0808 keV (the limit in XSPEC) and fitting a cooling flow model, we recover a much lower mass deposition rate of M⊙ yr-1, in line with the results of T. E. Clarke et al. (2004a). If we instead fit for the cooling within a smaller radius of 40″[58 kpc] encompassing the cD galaxy, we find M⊙ yr-1 which, again, is still higher than the observed SFR of the cD galaxy.
@@
Comparison to the simulations of J. A. ZuHone et al. (2018) suggests a scenario of a 1:10 mass ratio off-axis merger, with an impact parameter of 500 kpc. In this simulation, a sloshing spiral is set off after the initial core passage of the perturbing subsystem. Upon the second core passage, the subsystem drives a shock front and leaves behind a wake of material, similar to the “splash” observed in A2029. The particular epoch that matched the A2029 observations corresponds to Gyr after the start of the merger simulation (i.e., ). The agreement between simulation and observation supports a scenario in which A2029 is currently experiencing the late-time aftermath of a minor merger, with sloshing motions, a NW shock front, and a displaced splash-like structure to the SE. These features reinforce the view that A2029 appears to have experienced a minor merger 4 Gyr ago. Even this object, which has been described as one of the most relaxed clusters in the universe (A. Dressler, 1979; D. A. Buote & J. C. Tsai, 1996; B. T. Dullo et al., 2017), shows evidence of its merger history after deep observations and detailed study.
@@ A.1 Focal Plane Temperature & Gain Effects on Spectral Fit Parameters
Spectra were extracted from an annular region, covering a roughly isothermal area, centered on the peak of the cluster emission (as found by the 2D -model fitting in §3.1.1), defined with outer radius and inner radius equal to the cool core radius plus 15″, (using the reported cool core radius of A2029 (C. L. Sarazin & R. W. O’Connell, 1993) and similar methods of G. Schellenberger et al. (2015)). For each FP bin range (i.e., considering only the data within each set of dashed lines in Fig. 21), the extracted spectra were fit with an absorbed APEC model (R. K. Smith et al., 2001), keeping the hydrogen column density — — and redshift — — frozen (see §4). Since the observations were all taken around the same time of year each year and the temperature dependent gain calibration is expected to vary from year to year, the new data set was divided into two groups: one for 2022 observations, and one for 2023 observations — i.e, fit parameters were tied within each year group, but not across year groups. For each year group, within a given FP bin range, we obtained the temperature (kT), abundance (Z), and normalization with and without including a fit on the gain slope and offset parameters. Spectra of the pre-2005 set were extracted from the same region but without applying any additional filtering. The pre-2005 set was fit as a group with no gain fitting model in Xspec.

### v2

@@ 4.2 Temperature and Abundance Maps
Spectral maps of A2029 were created following the methods of Randall2008; Randall2009. For each spectral map pixel (98 pixel-1), spectra were extracted from circular regions whose radius was allowed to increase until a minimum of 20,000 background-subtracted counts were contained. The radius of the extraction region varies from 10″ ( 14 kpc) in bright regions near the core to ( 509 kpc) in faint regions near the edges of the map. Within this region, the spectrum of each ObsID is extracted and then fit simultaneously using an absorbed APEC model, with only the hydrogen column density — — and redshift — — frozen during fitting.
@@
The innermost spiral excess (highlighted by the white contours) traces keV gas temperatures and higher metallicities, as expected from sloshing ICM. The SW “splash” feature (white arrows in Fig. 7) traces regions of keV gas and leads into the coldest regions, with temperatures around keV. This cold region was previously seen in XMM and Chandra observations of A2029 (Bourdin2008; Walker2012). The splash also leads into a region of higher ( Z⊙) metallicity (Fig. 8). The region of isolated excess emission to the SW of the cluster (green arrow in Fig. 7) corresponds to a region of colder ( keV) gas, which may be associated with a background group or cluster (discussed further in §5.5). A large ( 0.9 Mpc) hot ( keV) region to the NW of the cluster center may indicate shock-heated gas in the region.
@@ 5.2 Potential Merger Shock
We used the python package pyproffit (v0.7.0; Eckert2020) to plot the surface brightness profile in a region over the large hot region observed to the NW in the temperature map (Fig. 7). We used a background subtracted image with point sources removed, an exposure map in units of seconds, and a user-generated error map to extract the surface brightness profile of the shock region using pyproffit. The surface brightness profile was extracted in a circular sector spanning 1.5∘ to , centered on the peak of the X-ray emission, with a maximum radial extent of 630″, and a bin size of 10″ in order to ensure a minimum of 5,000 net counts in the outermost bin.
We fit the 2D surface brightness within the kpc [300″] to kpc [630″] range using a projected broken power-law 3D density model (Markevitch2007; Owers2009; Tiwari2022; Rajpurohit2024) of the form:
@@
We can determine the Mach number of the potential shock front using a similar equation as Eq. 2, but using the temperature jump, expressed as a function of the density jump (Sanders2016a):
where and are the upstream and downstream gas temperatures, respectively. For regions where , when the temperature ratio, , is greater than unity it suggests the presence of a shock front, while ratios of less than unity indicate a potential cold front (Rajpurohit2024).
@@ 5.3.1 Bay Feature and SW Excess
As noted in §3, there appears to be a concave bay-like structure located kpc south of the cluster core that could be attributed to a Kelvin-Helmholtz instability (KHI). Similar structures have been reported in previous clusters (e.g., Perseus, Centaurus, and Abell 1795; Sanders2016; Walker2014; Walker2017) and are thought to be associated with large-scale KHI bays.
@@ 5.3.2 SE ‘Splash’
In §3, we noted a region of excess emission SE of the cluster core, which we dubbed the “splash” feature, noted in Fig. 3 with white arrows. Figure 15 illustrates residual emission (left panel) overlaid with the annular sectors used to extract the azimuthal surface brightness profiles (right panel). We compared the SE direction (solid squares) with the northwestern (NW) direction (open circles) in the right panel. The SW splash excess is observed between 196∘ to 247∘, highlighted in purple in Fig. 15. The source of this extended feature could be due to the second, perturbing cluster system. Some simulations (ZuHone2010; ZuHone2016) have shown similar “splash”-like features, in conjunction with sloshing spiral structures, that results after the second core passage of the perturbing body.
@@ 5.3.3 AGN Bubbles and Feedback
At the core of A2029 lies the AGN radio source PKS1508+059. AGN are known to inflate bubbles in the ICM that can offset radiative cooling. Previous studies of A2029 have been unable to detect any significant555Rafferty2006 report the detection of one cavity, but at low significance. evidence of such bubbles in X-ray imaging. While the observations here similarly reveal no clear evidence for such bubbles, we also cannot rule out the possibility that the southern bay feature could be associated with the inner edge of an outer ghost cavity that could have been inflated from the southern lobe. A similar situation arose in the Ophiuchus cluster, where a concave X-ray edge was initially deemed unrelated to the radio emission, but much deeper low-frequency radio data later revealed a matching ghost cavity (Giacintucci2020).
Figure 16 shows the X-ray and residual emission images in the inner 5′ 5′ cluster core to highlight the central WAT source. The overlaid cyan contours are the 1490 MHz radio emission taken from VLA observations presented in Taylor1994. Contours are drawn at levels corresponding to 0.0006, 0.0024, 0.0079, and 0.017 Jy beam-1. The X-ray images show no obvious signs of isolated regions of decreased emission near the radio lobes, nor anywhere in the ICM at larger radii. In the -model subtracted residual emission image, though showing decreased emission in the regions of both lobes, the observed depressions appear to be more closely associated with the cooling spiral structure rather than with distinct cavities. However, cavities may be particularly hard to detect in these regions because of confusion with the sloshing features.
@@
Recent work by Timmerman2022 applied a hybrid X-ray–radio method to estimate AGN cavity powers in galaxy clusters. For A2029, they utilize high-resolution () observations from the International Low Frequency Array (LOFAR; VanHaarlem2013) telescope (ILT) and low-resolution imaging from the LOFAR Two-Metre Sky Survey (LoTSS; Shimwell2017; Shimwell2019; Shimwell2022). While the authors report estimates of cavity volumes and power, they do not identify definitive X-ray cavities coincident with the radio lobes, noting that the complex spiral morphology of the ICM complicates the clear identification of such features.
We found in §4 a mass deposition rate within the central 40″ (equivalent to 58 kpc at the redshift of A2029) of M⊙ yr-1 when the low temperature of the cooling flow model is free. However, we find when the low temperature is fixed to 0.0808 keV and the high temperature is tied to an APEC component. The latter is the more relevant quantity to compare to the star formation rate of the cD galaxy — Hicks2010 measure a SFR of only M⊙ yr-1 and Fraser-McKelvie2014 estimate M⊙ yr-1. The large discrepancy in the results between the free- mass deposition rate and the SFR, coupled with the much smaller fixed- mass deposition rate, indicates that in A2029, most gas does not cool to sub-keV temperatures, consistent with a hidden or partial cooling flow scenario (Fabian2022).
To see if the central AGN is providing the heating necessary to offset the observed cooling, we compared the energy injection rate, , from the radio source to the luminosity of the cooling gas, . For A2029, PM2013 estimated an energy injection rate666Calculated as , where and are the cavity pressure and volume, respectively, averaged over a typical AGN repetition rate of yr. of erg s-1. Additionally, Timmerman2022 used low-resolution LOFAR observations to estimate a total cavity power (i.e., the AGN injection rate) of erg s-1; however the authors assumed radial propagation of the radio lobes, while it’s clear that for A2029, the lobes are bent relative to the jet axis.
@@ 5.4 Comparison to Simulation
We compare our observations to simulations of galaxy cluster mergers in order to get a better picture of the past merger history that would result in some of the features we have identified in A2029 so far. We examined the idealized binary merger simulations from the Galaxy Cluster Merger Catalog777https://gcmc.hub.yt (ZuHone2018) and found that our observations were qualitatively reproduced by a simulation of a 1:10 mass ratio merger, with an impact parameter of 500 kpc, originally presented in Zuhone2011a.
@@ 5.5 Optical Structure
Sohn2019 identify spectroscopic members of A2029, which we show overlaid on the X-ray image in the left panel of Fig. 20 in green circles. We create contours (in steps of 0.28 galaxies arcmin-1) of galaxy surface density, following Sohn2019, which we show as white contours in all panels of Fig. 20.
@@
In addition to member galaxies, Sohn2019 show that there are two infalling subsystems, consisting of Abell 2033 and a newly identified group, the Southern Infalling Group (SIG), based on weak lensing and X-ray observations, and phase-space analysis of the member galaxies. Both systems lie outside the Chandra FoV, which is roughly the size of the cluster radius Mpc (Sohn2019).
In the X-ray and residual emission images, we identified a region of excess emission located kpc SW of the cluster core (highlighted with green arrows in Figs. 3; see §3.1.1). In §5.3, we determined that this emission was not likely to be associated with the main cluster and could instead suggest the presence of a foreground or background group or cluster, with its own intragroup or intracluster medium that we are detecting. The thumbnail image of the radio source, shown in Fig. 2 panel (b), reveals the presence of numerous red sources clustered together. Sohn2019 measures a spectroscopic redshift of , attributed to the largest and brightest source of the group shown in Fig. 2 panel (b). The concurrence with radio emission indicates the presence of an AGN within the group of red galaxies. In addition, this group coincides with the region of excess X-ray emission seen to the SW of the cluster (green arrow in Fig. 3). All this together supports the picture of a background group or cluster, with its own intragroup or intracluster medium, that lies at a redshift .
The southernmost double-lobe radio source, located 680 kpc nearly due south of the cluster center, is at a redshift of (Sohn2019), indicating it could be associated with the SW radio source, which lies at similar redshift. The thumbnail of the source, shown in panel (c) of Fig. 2, shows two sources near the expected location of the radio source, with the one identified as a point source during the wavdetect run (see §3) indicated by a red circle. This source is likely the AGN driving the two radio lobes seen in 1.4 GHz.
@@ 6 Summary and Conclusions
We find a global, best-fit projected temperature of keV and abundance of within the inner 116″[ kpc]. This is slightly lower temperature than previous studies (e.g., PM2013) but consistent within the combined statistical and systematic uncertainties. A two-temperature model provides a significantly improved fit, with components at keV and keV, yielding much tighter constraints on the hot temperature component than earlier studies (Clarke2004; PM2013). Fitting a cooling flow model returned best-fit temperatures of keV and keV, abundance , and mass deposition rate M⊙ yr-1. This is in agreement with previous studies that have reported similarly high cooling flow rates (Sarazin1992; Peres1998; PM2013), despite the relatively low levels of star formation in the central cD galaxy compared to expectations from unimpeded cooling (McNamara2007; McDonald2010; Hicks2010). The low temperature component is consistent with the latest XRISM findings which report a 3.42 keV temperature component detected within the inner kpc [2.5′] (Sarkar2025). However, when keeping the low temperature fixed to 0.0808 keV (the limit in XSPEC) and fitting a cooling flow model, we recover a much lower mass deposition rate of M⊙ yr-1, in line with the results of Clarke2004. If we instead fit for the cooling within a smaller radius of 40″[58 kpc] encompassing the cD galaxy, we find M⊙ yr-1 which, again, is still higher than the observed SFR of the cD galaxy.
@@
Comparison to the simulations of ZuHone2018 suggests a scenario of a 1:10 mass ratio off-axis merger, with an impact parameter of 500 kpc. In this simulation, a sloshing spiral is set off after the initial core passage of the perturbing subsystem. Upon the second core passage, the subsystem drives a shock front and leaves behind a wake of material, similar to the “splash” observed in A2029. The particular epoch that matched the A2029 observations corresponds to Gyr after the start of the merger simulation (i.e., ). The agreement between simulation and observation supports a scenario in which A2029 is currently experiencing the late-time aftermath of a minor merger, with sloshing motions, a NW shock front, and a displaced splash-like structure to the SE. These features reinforce the view that A2029 appears to have experienced a minor merger 4 Gyr ago. Even this object, which has been described as one of the most relaxed clusters in the universe (Dressler1979; Buote1996; Dullo2017), shows evidence of its merger history after deep observations and detailed study.
@@ A.1 Focal Plane Temperature & Gain Effects on Spectral Fit Parameters
Spectra were extracted from an annular region, covering a roughly isothermal area, centered on the peak of the cluster emission (as found by the 2D -model fitting in §3.1.1), defined with outer radius and inner radius equal to the cool core radius plus 15″, (using the reported cool core radius of A2029 (Sarazin1992) and similar methods of Schellenberger2015). For each FP bin range (i.e., considering only the data within each set of dashed lines in Fig. 21), the extracted spectra were fit with an absorbed APEC model (Smith2001), keeping the hydrogen column density — — and redshift — — frozen (see §4). Since the observations were all taken around the same time of year each year and the temperature dependent gain calibration is expected to vary from year to year, the new data set was divided into two groups: one for 2022 observations, and one for 2023 observations — i.e, fit parameters were tied within each year group, but not across year groups. For each year group, within a given FP bin range, we obtained the temperature (kT), abundance (Z), and normalization with and without including a fit on the gain slope and offset parameters. Spectra of the pre-2005 set were extracted from the same region but without applying any additional filtering. The pre-2005 set was fit as a group with no gain fitting model in Xspec.